# Large Language Models:

### v1

@@ 7 Conclusion
Machine learning tools are radically expanding the scope of empirical research in economics. We now move beyond estimating average causal effects to learning personalized treatment effects (e.g., athey2017econometrics; wager2018estimation). We use unstructured data, such as satellite images, to infer outcomes at high-frequencies and granular scales (e.g., DonaldsonStoreygard(16); rambachan2024programevaluationremotelysensed). We tackle prediction policy problems (kleinberg2015prediction; kleinberg2018human) and develop tools for hypothesis generation (fudenberg2019predicting; ludwig2024machine; mullainathan2024predictive). In this context, LLMs are the latest machine learning tools to enter our empirical toolkit.
While powerful and easy to use, LLMs are also brittle, which has consequences for empirical research using LLM outputs. As we showed, naive uses of LLM outputs distort both downstream predictions and parameter estimates. To understand this problem and identify constructive solutions, we have developed here an econometric framework for studying LLMs. Rather than making assumptions about how they are designed, we treat LLMs as black boxes and provided conditions that the LLM’s outputs must satisfy in order to be used in prediction and estimation problems.
The need for such econometric contracts pervades every aspect of our efforts to incorporate machine learning and other modern computing tools into economics research. While such contracts are familiar to empirical researchers, their development will need to be done differently for this new world of machine learning. Traditionally, econometrics provided contracts for empirical tools that we designed ourselves. Today, economists increasingly sit on the consumer side of new AI models, and we are blind to most of what happens in their production. As we have argued in this paper (using LLMs as an example), it is nonetheless still possible to write meaningful contracts for AI tools that enable economists to use these tools to push the empirical frontier rigorously forward.
@@ Appendix A Appendix Figures and Tables
Distance Average Benchmark
@@
Distance Average Benchmark
Cosine similarity 0.829 0.379
Euclidean distance 0.538 1.110
@@
[…]
Notes: This table calculates the cosine similarity and Euclidean distance between embeddings of GPT-4o’s completed bill descriptions and embeddings of the original bill descriptions. We construct embeddings using OpenAI’s text-embedding-3-small model. As a benchmark, we calculate the average cosine similarity and Euclidean distance between 10,000 randomly selected pairs of original bill descriptions. Table (a) provides results for the base prompt and Table (b) provides results for the base prompt with the additional date restriction. The results in Table (a) and Table (b) are the same up to 3 decimal places. See Section 4.3.1 for discussion.
@@
Distance Average Benchmark
@@
Distance Average Benchmark
@@
[…]
Notes: This table calculates the cosine similarity and Euclidean distance between embeddings of GPT-4o’s completed financial news headlines and embeddings of the original financial news headlines. We construct embeddings using OpenAI’s text-embedding-3-small model. As a benchmark, we calculate the average cosine similarity and Euclidean distance between 10,000 randomly selected pairs of original financial news headlines. Table (a) provides results for the base prompt and Table (b) provides results for the base prompt with the additional date restriction. The results in Table (a) and Table (b) are the same up to 3 decimal places. See Section 4.3.2 for discussion.
Return Horizon
1 day 5 day 10 day
Point Estimates
Mean -2.18 -1.24 -0.05
Median -2.3 -1.13 0.03
5th Percentile -5.16 -2.86 -2.29
95th Percentile 0.46 -0.02 2.31
Sample Average 0.05 0.31 0.58
(a) Positive Labels
Return Horizon
1 day 5 day 10 day
Point Estimates
Mean 2.57 0.85 0.41
Median 2.78 2.00 2.38
5th Percentile -1.31 -6.73 -9.39
95th Percentile 5.55 4.52 5.85
Sample Average 0.05 0.31 0.58
(b) Negative Labels
Table A3: Point estimates across large language models and prompting strategies on financial news headlines using confidence.
[…]
Notes: On financial news headlines from 2019, we prompt GPT-3.5-Turbo, GPT-4o-mini, and GPT-4o to label each headline for whether it expressed positive, negative or uncertain news about the respective company using alternative prompting strategies. For each model $m$ and prompt $p$, we regress the realized returns of each stock within 1-day of the headline’s publication date on each large language model’s labels $\widehat{V}_{r}^{m,p}$, the large language model’s assessed confidence denoted $C_{r}^{m,p}$ and their interaction, controlling for lagged realized returns. See Section 5.3.2 for discussion.
Original Bill: a bill to authorize the federal programs to prevent violence against women and for other purposes GPT-4o: a bill to authorize the federal programs to prevent violence against women and for other purposes
Original Bill: a bill to repeal the public utility holding company act of 1935 to enact the public utility holding company act of 1997 and for other purposes GPT-4o: a bill to repeal the public utility holding company act of 1935 to enact the public utility holding company act of 1997 and for other purposes
@@
Figure A1: Examples of GPT-4o completions with date restriction that exactly match original descriptions of Congressional legislation.
[…]
Notes: On 10,000 randomly sampled Congressional bills, we prompted GPT-4o to complete the description of the Congressional bill based on 50% of its text. The prompt included an explicit date restriction. See Section 4.3.1 for discussion.
Original Headline: report comcast might sell hulu stake to disney GPT-4o: report comcast might sell hulu stake to disney
Original Headline: euronet worldwide shares are trading lower after da davidson downgraded the stock from buy to neutral GPT-4o: euronet worldwide shares are trading lower after da davidson downgraded the stock from buy to neutral
@@
Figure A2: Examples of GPT-4o completions with date restriction that exactly match original financial news headlines.
[…]
Notes: On 10,000 randomly sampled financial news headlines from 2019, we prompted GPT-4o to complete the financial news headline based on 50% of its text. The prompt included an explicit date restriction. See Section 4.3.2 for discussion.
@@
Figure A3: Variation in t-statistics realized returns across large language models and prompting strategies on financial news headlines using model’s reported confidence.
[…]
Notes: On financial news headlines from 2019, we prompt GPT-3.5-Turbo, GPT-4o-mini, and GPT-4o to label each headline for whether it expressed positive, negative or uncertain news about the respective company using alternative prompting strategies. For each model $m$ and prompt $p$, we regress the realized returns of each stock within 1 day, 5 days, and 10 days of the headline’s publication date on each large language model’s labels $\widehat{V}_{r}^{m,p}$, the large language model’s reported confidence denoted $C_{r}^{m,p}$ and their interaction, controlling for lagged realized returns. We separately report the t-statistics associated with the regression coefficients on whether the headline is labeled as positive or negative news (standard errors are two-way clustered at the date and company level). In each subplot, the t-statistics are sorted in ascending order for clarity. See Section 5.3.2 for discussion.
@@
[…]
Notes: On 10,000 randomly sampled Congressional bills, we prompt GPT-3.5-turbo and GPT-4o to label the policy topic of each Congressional bill. For each pair of prompting strategies, we calculate the fraction of congressional bills that receive the same label, separately by large language model. See Section 5.3.3 for discussion.
@@ Appendix B Proofs of Main Results
B.1 Proof of Lemma 1, Proposition 1, and Corollary 1
@@
$\mathbb{E}_{Q}[\sum_{r\in\mathcal{R}}D_{r}\ell(Y_{r},\widehat{m}(r;t))]-\mathbb{E}_{Q}[\sum_{r\in\mathcal{R}}D_{r}\ell(Y_{r},\widehat{m}(r;t))\mid T=t]=$
$\sum_{r\in\mathcal{R}}q_{r}^{D}\ell(Y_{r},\widehat{m}(r;t))-\sum_{r\in\mathcal{R}}q_{r}^{D\mid T}(t_{r})\ell(Y_{r},\widehat{m}(r;t))=\sum_{r\in\mathcal{R}}\left(q_{r}^{D}-q_{r}^{D\mid T}(t_{r})\right)\ell(Y_{r},\widehat{m}(r;t)).$
@@
$\mathbb{E}_{Q}[\sum_{r\in\mathcal{R}}D_{r}\ell(Y_{r},\widehat{m}(r;t))]-\mathbb{E}_{Q}[\sum_{r\in\mathcal{R}}D_{r}\ell(Y_{r},\widehat{m}(r;t))\mid T=t]=\sum_{r\in\mathcal{R}}q_{r}^{D}\left(1-\frac{q_{r}^{T\mid D}(t_{r})}{q_{r}^{T}(t_{r})}\right)\ell(Y_{r},\widehat{m}(r;t)).$
Lemma 1 then follows immediately.
Finally, to prove Corollary 1, suppose that $q_{r}^{D}\in\{0,1\}$ for all $r\in\mathcal{R}$. In this case,
$\sum_{r\in\mathcal{R}}q_{r}^{D}\left(1-\frac{q_{r}^{T\mid D}(t_{r})}{q_{r}^{T}(t_{r})}\right)\ell(Y_{r},\widehat{m}(r;t))=\sum_{r\colon q_{r}^{D}=1}\left(1-\frac{q_{r}^{T\mid D}(t_{r})}{q_{r}^{T}(t_{r})}\right)\ell(Y_{r},\widehat{m}(r;t)).$
Furthermore, for all $r$ such that $q_{r}^{D}=1$, we have $q_{r}^{T\mid D}(t_{r})=Q(T_{r}=t_{r},D_{r}=1)$ equals $q_{r}^{T}(t_{r})=Q(T_{r}=t_{r})$ by the law of total probability. Consequently, $01-\frac{q_{r}^{T\mid D}(t_{r})}{q_{r}^{T}(t_{r})}=$ 0 for all $r$ such that $q_{r}^{D}=1$ and no training leakage is satisfied. $\Box$
@@ B.2 Proof of Lemma 2
The result then follows by applying the same argument as the proof of Lemma 1 to rewrite the first term as $\mathbb{E}_{Q}[\sum_{r\in\mathcal{R}}D_{r}\left(\frac{q^{T\mid D}(t_{r})}{q_{r}^{T}(t_{r})}-1\right)g(\widehat{m}(r),W_{r};\theta)]$. $\Box$
@@ B.3 Proof of Lemma 3 and Proposition 2
$\left|\sum_{r\in\mathcal{R}}q_{t}^{D}\left(g(\widehat{m}(t),W_{r};\theta)-g(V_{r},W_{r};\theta)\right)\right|.$
@@ B.4 Proof of Proposition 3
To show (1), given that $q_{r}^{T}(t_{r})=q_{r}^{T\mid D}(t_{r})$ for all $r\in\mathcal{R}$, it follows that
$\beta=\left(\sum_{r\in\mathcal{R}}q_{r}^{D}W_{r}W_{r}^{\prime}\right)^{-1}\left(\sum_{r\in\mathcal{R}}q_{r}^{D}W_{r}\widehat{V}_{r}\right)\mbox{ and }\beta^{*}=\left(\sum_{r\in\mathcal{R}}q_{r}^{D}W_{r}W_{r}^{\prime}\right)^{-1}\left(\sum_{r\in\mathcal{R}}q_{r}^{D}W_{r}V_{r}\right).$
@@
To show (2), since there is again no training leakage by assumption, it follows that
$\beta=\left(\sum_{r\in\mathcal{R}}q_{r}^{D}\widehat{V}_{r}\widehat{V}_{r}^{\prime}\right)^{-1}\left(\sum_{r\in\mathcal{R}}q_{r}^{D}\widehat{V}_{r}W_{r}\right)\mbox{ and }\beta^{*}=\left(\sum_{r\in\mathcal{R}}q_{r}^{D}V_{r}V_{r}^{\prime}\right)^{-1}\left(\sum_{r\in\mathcal{R}}q_{r}^{D}V_{r}W_{r}\right).$
@@ C.1.1 The Researcher’s Sample Average Loss
To tackle the prediction problem, the researcher calculates the sample average loss of the large language model’s predictions on their collected dataset:
@@
[…]
@@ C.1.2 The Researcher’s Sample Moment Condition
[…]
@@
By the same argument, we can interpret the researcher sampling distribution over text pieces conditional on the large language model’s realized training dataset as i.n.i.d sampling from the finite population of text pieces; and the researcher’s plug-in moment function then calculates the sample mean of the finite population characteristics $g(\widehat{m}(r;t),W_{r};\theta)$. Existing results then provide regularity conditions under which Equation (6) holds and
[…]
@@ C.2 Analyzing the Asymptotic Distribution of Bias-Corrected Coefficient
In this section, we separately analyze the asymptotic distribution of the bias-corrected regression coefficient introduced in Section 5.4.1 in two separate cases: first, when the economic concept $V_{r}$ is the dependent variable; and second, when the economic concept $V_{r}$ is the independent variable.
@@ C.2.1 Linear Regression with Large Language Model Labels as the Dependent Variable
As discussed in Section 5.4.1 of the main text, we now study the limiting distribution of the bias-corrected linear regression in which the researcher uses the economic concept as the dependent variable. It is convenient to now define the researcher’s sampling indicator as taking three possible $D_{r}\in\{0,1,2\}$, where $D_{r}=0$ denotes the researcher does not sample the text piece $r$, $D_{r}=1$ denotes that the researcher samples the text piece in the primary sample and observes $(\widehat{m}(r;t),W_{r})$, and $D_{r}=2$ denotes that the researcher samples the text piece in the validation sample and observes $(\widehat{m}(r;t),V_{r},W_{r})$. Altogether the researcher observes $(\widehat{m}(r;t),W_{r})$ for all $r\in\mathcal{R}$ with $D_{r}=1$ and $(\widehat{m}(r;t),V_{r},W_{r})$ for all $r\in\mathcal{R}$ with $D_{r}=2$.
On the primary sample, the researcher calculate the plug-in regression coefficient
@@
To derive the limiting distribution as the number of economically relevant text pieces $N_{R}$ grows large, we make three simplifying assumptions. First, we assume that $W_{r}$ is a scalar, which is not technically necessary but will simplify the resulting expressions. Second, we assume the large language model satisfies no training leakage as mentioned in the the main text. Third, we analyze a research context $Q(\cdot)$ in which the researcher randomly samples text pieces into their dataset and further randomly partitions the collected text pieces into the primary and validation sample. More formally, the text pieces are randomly sampled into three groups of size $N_{o},N_{p},N_{v}$ respectively and the probability that the vector $D$ takes a particular value $d$ is given by $N_{o}!N_{p}!N_{v}!/N_{R}!$, where $d$ satisfies $\sum_{r\in\mathcal{R}}1\{D_{r}=0\}=N_{o}$, $\sum_{r\in\mathcal{R}}1\{D_{r}=1\}=N_{p}$, $\sum_{r\in\mathcal{R}}1\{D_{r}=2\}=N_{v}$. Finally, we will assume there exists some finite constant $M>0$ such that $-M\leq W_{r},V_{r},\widehat{m}(r;t)\leq M$ for all $r\in\mathcal{R}$. The last two assumptions enable us to apply existing finite-population central limit theorem in deriving limiting distributions
[…]
@@
[…]
@@
[…]
@@
and $\sigma^{2}_{W}$ the limit of $\frac{1}{N}\sum_{r\in\mathcal{R}}W_{r}^{2}$. This delivers Equation (13) given in Section 5.4.1 of the main text. By a similar argument, we can show that the validation-sample only regression coefficient has a limiting distribution given by
[…]
for $\Omega^{validation}=\sigma^{-4}_{W}\frac{1-\rho_{V}}{\rho_{V}}\sigma^{2}_{WV}$, as stated in Section 5.4.1 of the main text.

### v3

@@ 7 Conclusion
Machine learning and artificial intelligence expand the scope of empirical research in economics. We now move beyond estimating average causal effects to learning personalized treatment effects (e.g., athey2017econometrics; wager2018estimation). We use unstructured data, such as satellite images and digital traces, to infer outcomes at high-frequencies and granular scales (e.g., DonaldsonStoreygard(16); BlumenstockEtAl(15); rambachan2024programevaluationremotelysensed). We tackle prediction policy problems (kleinberg2015prediction; kleinberg2018human; Mullainathan(25)) and develop algorithms for hypothesis generation (fudenberg2019predicting; ludwig2024machine; mullainathan2024predictive). LLMs are the latest tools to enter empirical work.
By radically reducing the cost of analyzing vast text corpora, LLMs enable economists to tackle questions previously impossible due to scale or expense. Using large language models, researchers can predict market reactions from earnings calls, measure sentiment across historical newspapers, track partisan polarization in social media, and simulate human responses at minimal cost. Yet these are complex algorithms that seemingly resist traditional econometric analysis. Our framework shows how to harness LLMs despite their complexity. The straightforward practices we recommend unlock LLMs’ transformative potential for empirical research.
Main Figures and Tables
Figure 1: Variation in t-statistics for realized returns across large language models and prompting strategies on financial news headlines.
Figure 2: Variation in t-statistics across large language models and prompting strategies on congressional legislation.
Figure 3: Normalized bias of the plug-in regression and bias-corrected regression across Monte Carlo simulations based on congressional legislation.
Figure 4: Cumulative distribution function of mean square error for the bias-corrected estimator against validation-sample only estimator.
Return Horizon
Point Estimates 1 day 5 days 10 days
Mean -0.677 -0.290 0.289
Median -0.221 0.141 0.637
$5^{\text{th}}$ Percentile -2.475 -1.787 -1.584
$95^{\text{th}}$ Percentile 0.079 0.465 1.186
Sample Average 0.045 0.316 0.588
(a) Positive Labels
Return Horizon
Point Estimates 1 day 5 days 10 days
Mean 1.219 1.027 1.088
Median 0.862 1.082 1.156
$5^{\text{th}}$ Percentile 0.092 -0.201 -0.126
$95^{\text{th}}$ Percentile 3.474 2.325 2.505
Sample Average 0.045 0.316 0.588
(b) Negative Labels
Table 1: Variation in point estimates across large language models and prompting strategies on financial news headlines.
Policy Topic Covariate Point Estimates Sample
Mean Median 5% 95% Average
Health DW1 -0.018 -0.020 -0.023 -0.006 0.150
Health Democrat 0.012 0.014 0.003 0.016 0.150
Health Senate 0.005 0.004 0.001 0.012 0.150
Banking, Finance & Domestic Com. DW1 0.013 0.013 0.007 0.022 0.127
Banking, Finance & Domestic Com. Democrat -0.010 -0.010 -0.015 -0.004 0.127
Banking, Finance & Domestic Com. Senate -0.008 -0.009 -0.016 -0.001 0.127
Defense DW1 0.022 0.022 0.015 0.034 0.204
Defense Democrat -0.005 -0.004 -0.009 0.000 0.204
Defense Senate -0.008 -0.006 -0.019 -0.003 0.204
Government Operations DW1 0.015 0.015 0.000 0.028 0.281
Government Operations Democrat -0.005 -0.004 -0.014 0.003 0.281
Government Operations Senate -0.013 -0.013 -0.020 -0.006 0.281
Public Lands & Water Management DW1 0.027 0.027 0.020 0.030 0.238
Public Lands & Water Management Democrat -0.006 -0.007 -0.010 -0.003 0.238
Public Lands & Water Management Senate 0.032 0.032 0.025 0.036 0.238
Table 2: Variation in point estimates across large language models and prompting strategies on Congressional bills.
Median 5% 95%
Normalized Bias
Plug-In -0.023 -1.899 2.211
Debiased 0.001 -0.055 0.066
Coverage
Plug-In 0.820 0.381 0.945
Debiased 0.930 0.910 0.945
(a) GPT-3.5-turbo
Median 5% 95%
Normalized Bias
Plug-In 0.084 -1.411 1.514
Debiased 0.001 -0.055 0.054
Coverage
Plug-In 0.920 0.637 0.954
Debiased 0.927 0.902 0.945
(b) GPT-4o
Median 5% 95%
Normalized Bias
Plug-In 0.030 -1.646 1.567
Debiased -0.002 -0.052 0.054
Coverage
Plug-In 0.906 0.589 0.953
Debiased 0.927 0.903 0.945
(c) GPT-5-mini
Median 5% 95%
Normalized Bias
Plug-In 0.168 -2.079 2.092
Debiased 0.005 -0.052 0.060
Coverage
Plug-In 0.779 0.387 0.953
Debiased 0.930 0.906 0.946
(d) GPT-5-nano
Table 3: Summary statistics for normalized bias and coverage across Monte Carlo simulations based on Congressional legislation.
Large Language Models:
An Applied Econometric Perspective
Online Appendix
Jens Ludwig & Sendhil Mullainathan & Ashesh Rambachan
@@ Appendix A Appendix Figures and Tables
Table A1: Accuracy, true positive rate (TPR), and false positive rate (FPR) of GPT-4o’s predictions on Congressional legislation.
Accuracy TPR FPR
House 0.912 0.198 0.031
Senate 0.925 0.225 0.031
(a) Base prompt
Accuracy TPR FPR
House 0.644 0.691 0.359
Senate 0.695 0.711 0.306
(b) Prompt with date restriction
@@
Metric Average Benchmark
@@
Metric Average Benchmark
Cosine similarity 0.830 0.379
Euclidean distance 0.536 1.110
@@
Metric Average Benchmark
@@
Metric Average Benchmark
@@
Original Bill: to amend title xviii of the social security act to distribute additional information to medicare beneficiaries to prevent health care fraud and for other purposes GPT-4o: to amend title xviii of the social security act to distribute additional information to medicare beneficiaries to prevent health care fraud and for other purposes
Original Bill: a bill to amend the comprehensive environmental response compensation and liability act of 1980 to promote the cleanup and reuse of brownfields to provide financial assistance for brownfields revitalization to enhance state response programs and for other purposes GPT-4o: a bill to amend the comprehensive environmental response compensation and liability act of 1980 to promote the cleanup and reuse of brownfields to provide financial assistance for brownfields revitalization to enhance state response programs and for other purposes
@@
Figure A1: Two examples of GPT-4o completions that exactly match original descriptions of congressional legislation.
Original Headline: piper jaffray maintains overweight on activision blizzard raises price target to 62 GPT-4o: piper jaffray maintains overweight on activision blizzard raises price target to 62
Original Headline: sinclair completes acquisition of regional sports networks from disney GPT-4o: sinclair completes acquisition of regional sports networks from disney
@@
Figure A2: Two examples of GPT-4o completions that exactly match original financial news headlines.
@@
Figure A3: Variation in pairwise agreement between large language model labels across prompting strategies on financial news headlines.
@@
Figure A5: Accuracy of large language model labels of bill topic across model and prompt variation.
@@ Appendix B Proofs of Main Results
B.1 Proof of Lemma 1 and Proposition 1
@@
$\mathbb{E}_{Q}[\sum_{r\in\mathcal{R}}D_{r}\ell(Y_{r},\widehat{m}(r))]-\mathbb{E}_{Q}[\sum_{r\in\mathcal{R}}D_{r}\ell(Y_{r},\widehat{m}(r))\mid T=t]=$
$\sum_{r\in\mathcal{R}}q^{D}_{r}\ell(Y_{r},\widehat{m}(r))-\sum_{r\in\mathcal{R}}q_{r}^{D\mid T}(t_{r})\ell(Y_{r},\widehat{m}(r))=\sum_{r\in\mathcal{R}}\left(q_{r}^{D}-q_{r}^{D\mid T}(t_{r})\right)\ell(Y_{r},\widehat{m}(r)).$
@@
$\mathbb{E}_{Q}[\sum_{r\in\mathcal{R}}D_{r}\ell(Y_{r},\widehat{m}(r))]-\mathbb{E}_{Q}[\sum_{r\in\mathcal{R}}D_{r}\ell(Y_{r},\widehat{m}(r))\mid T=t]=\sum_{r\in\mathcal{R}}q_{r}^{D}\left(1-\frac{q_{r}^{T\mid D}(t_{r})}{q_{r}^{T}(t_{r})}\right)\ell(Y_{r},\widehat{m}(r)).$
Lemma 1 then follows immediately. $\Box$
@@ B.2 Proof of Lemma 2
The result then follows by applying the same argument as the proof of Lemma 1 to rewrite the first term as $\mathbb{E}_{Q}[\sum_{r\in\mathcal{R}}D_{r}\left(\frac{q^{T\mid D}_{r}(t_{r})}{q_{r}^{T}(t_{r})}-1\right)g(\widehat{m}(r),W_{r};\theta)]$. $\Box$
@@ B.3 Proof of Lemma 3 and Proposition 2
$\left|\sum_{r\in\mathcal{R}}q_{t}^{D}\left(g(\widehat{m}(r),W_{r};\theta)-g(V_{r},W_{r};\theta)\right)\right|.$
@@ B.4 Proof of Proposition 3
To show (i), given that $q_{r}^{T}(t_{r})=q_{r}^{T\mid D}(t_{r})$ for all $r\in\mathcal{R}$, it follows that
[…]
@@
To show (ii), since there is again no training leakage by assumption, it follows that
[…]
@@ C.1.1 The Researcher’s Sample Average Loss
To tackle the prediction problem, the researcher calculates the sample average loss of the large language model’s predictions:
@@
$\frac{1}{N}\sum_{r\in\mathcal{R}}D_{r}\ell(Y_{r};\widehat{m}(r;t))-\frac{1}{\mathbb{E}_{Q}\left[\sum_{r\in\mathcal{R}}D_{r}\mid T=t\right]}\mathbb{E}_{Q}\left[\sum_{r\in\mathcal{R}}D_{r}\ell(Y_{r};\widehat{m}(r;t))\mid T=t\right]\xrightarrow{p}0,$
@@ C.1.2 The Researcher’s Sample Moment Condition
$\frac{1}{N}\sum_{r\in\mathcal{R}}D_{r}g(W_{r},V_{r};\theta)-\frac{1}{\mathbb{E}_{Q}\left[\sum_{r\in\mathcal{R}}D_{r}\right]}\mathbb{E}_{Q}\left[\sum_{r\in\mathcal{R}}D_{r}g(W_{r},V_{r};\theta)\right]\xrightarrow{p}0,$
@@
By the same argument, we can interpret the researcher sampling distribution over text pieces conditional on the large language model’s realized training dataset as i.n.i.d sampling from the finite population of text pieces; and the researcher’s plug-in moment function then calculates the sample mean of the finite population characteristics $g(\widehat{m}(r;t),W_{r};\theta)$. Existing results then provide regularity conditions under which
$\frac{1}{N}\sum_{r\in\mathcal{R}}D_{r}g(\widehat{m}(r;t),W_{r};\theta)-\frac{1}{\mathbb{E}\left[\sum_{r\in\mathcal{R}}D_{r}\mid T=t\right]}\mathbb{E}\left[\sum_{r\in\mathcal{R}}D_{r}g(\widehat{m}(r;t),W_{r};\theta)\mid T=t\right]\xrightarrow{p}0$
@@ C.2 Analyzing the Asymptotic Distribution of Bias-Corrected Coefficient
In this section, we separately analyze the asymptotic distribution of the bias-corrected regression coefficient introduced in Section 4.3 in two separate cases: first, when the economic concept $V_{r}$ is the dependent variable; and second, when the economic concept $V_{r}$ is the independent variable.
@@ C.2.1 Linear Regression with Large Language Model Labels as the Dependent Variable
As discussed in Section 4.3, we study the limiting distribution of the bias-corrected linear regression in which the researcher uses the economic concept as the dependent variable. It is convenient to now define the researcher’s sampling indicator as taking three possible $D_{r}\in\{0,1,2\}$, where $D_{r}=0$ denotes the researcher does not sample the text piece $r$, $D_{r}=1$ denotes that the researcher samples the text piece in the primary sample and observes $(\widehat{m}(r;t),W_{r})$, and $D_{r}=2$ denotes that the researcher samples the text piece in the validation sample and observes $(\widehat{m}(r;t),V_{r},W_{r})$. Altogether the researcher observes $(\widehat{m}(r;t),W_{r})$ for all $r\in\mathcal{R}$ with $D_{r}=1$ and $(\widehat{m}(r;t),V_{r},W_{r})$ for all $r\in\mathcal{R}$ with $D_{r}=2$.
On the primary sample, the researcher calculates the plug-in regression coefficient
@@
To derive the limiting distribution as the number of economically relevant text pieces $N_{R}$ grows large, we make three simplifying assumptions. First, we assume that $W_{r}$ is a scalar, which is not technically necessary but will simplify the resulting expressions. Second, we assume the large language model satisfies no training leakage as mentioned in the the main text. Third, we analyze a research context $Q(\cdot)$ in which the researcher randomly samples text pieces into their dataset and further randomly partitions the collected text pieces into the primary and validation sample. More formally, the text pieces are randomly sampled into three groups of size $N_{o},N_{p},N_{v}$ respectively and the probability that the vector $D$ takes a particular value $d$ is given by $N_{o}!N_{p}!N_{v}!/N_{R}!$, where $d$ satisfies $\sum_{r\in\mathcal{R}}1\{D_{r}=0\}=N_{o}$, $\sum_{r\in\mathcal{R}}1\{D_{r}=1\}=N_{p}$, $\sum_{r\in\mathcal{R}}1\{D_{r}=2\}=N_{v}$. Finally, we will assume there exists some finite constant $M>0$ such that $M\leq W_{r},V_{r},\widehat{m}(r;t)\leq M$ for all $r\in\mathcal{R}$. The last two assumptions enable us to apply existing finite-population central limit theorem in deriving limiting distributions
We study the properties of the bias-corrected regression and the validation-sample only regression along a sequence of finite populations satisfying $N_{R}\rightarrow\infty$, $N_{v}/N_{R}=\rho_{v}>0$, $N_{p}/N_{R}=\rho_{p}>0$. Under these stated conditions, results in li_general_2017 imply that $\frac{1}{N_{p}}\sum_{r\in\mathcal{R}}1\{D_{r}=1\}W_{r}^{2}-\frac{1}{N_{R}}\sum_{r\in\mathcal{R}}W_{r}^{2}\xrightarrow{p}0$ and $\frac{1}{N_{v}}\sum_{r\in\mathcal{R}}1\{D_{r}=2\}W_{r}^{2}-\frac{1}{N_{R}}\sum_{r\in\mathcal{R}}W_{r}^{2}\xrightarrow{p}0$. We therefore focus on analyzing the properties of $\frac{1}{N_{p}}\sum_{r\in\mathcal{R}}1\{D_{r}=1\}W_{r}\widehat{m}(r;t)$ and $\frac{1}{N_{v}}\sum_{r\in\mathcal{R}}1\{D_{r}=1\}W_{r}\Delta_{r}$.
@@
$\sqrt{N_{R}}\left((\bar{X}_{p},\bar{Z}_{v})^{\prime}-(\bar{X}_{N},\bar{Z}_{N})^{\prime}\right)\xrightarrow{d}N\left(0,\begin{pmatrix}\frac{1-\rho_{p}}{\rho_{p}}\sigma^{2}_{X}&-\sigma_{X}\sigma_{Z}\\ -\sigma_{X}\sigma_{Z}&\frac{1-\rho_{v}}{\rho_{v}}\sigma^{2}_{Z}\end{pmatrix}\right)$.
@@
$\sqrt{N}_{R}\left(\widehat{\beta}^{debiased}-\beta^{*}\right)\xrightarrow{d}N(0,\Omega^{debiased})$
@@
and $\sigma^{2}_{W}$ the limit of $\frac{1}{N}\sum_{r\in\mathcal{R}}W_{r}^{2}$. This delivers Equation (15) given in Section 4.3 of the main text. By a similar argument, we can show that the validation-sample only regression coefficient has a limiting distribution given by
$\sqrt{N}_{R}\left(\widehat{\beta}^{validation}-\beta^{*}\right)\xrightarrow{d}N(0,\Omega^{validation})$
for $\Omega^{validation}=\sigma^{-4}_{W}\frac{1-\rho_{V}}{\rho_{V}}\sigma^{2}_{WV}$, as stated in Section 4.3 of the main text.